Question 4

Blue Distribution markets CDs of the performing artist Marilynn. At the beginning of October, Blue had in beginning inventory 4,420 of Marilynn’s CDs with a unit cost of £2. During October, Blue made the following purchases of Marilynn’s CDs.

Oct. 3 6,630 @ £6 Oct. 19 8,840 @ £13
Oct. 9 12,155 @ £7 Oct. 25 4,420 @ £14


During October, 29,835 units were sold. Blue uses a periodic inventory system.

Determine (1) the ending inventory and (2) the cost of goods sold under the two assumed cost flow methods (FIFO and average-cost). (Round average unit cost to 4 decimal places, e.g. 25.1672 and final answers to 0 decimal places, e.g. 2,120.)

FIFO

AVERAGE-COST

The ending inventory £

£

The cost of goods sold £

£
Answer #1

The answer has been presented in the supporting sheet. All the parts has been solved with detailed explanation and format. For detailed answer refer to the supporting sheet.

Answer 3 Firstly we have to calculate the cost of goods available for sale 4 = beginning inventory + purchases 5 = (4420*2)+(

4 Cost of goods sold fifo == total cost of goods available for sale - ending inventory FIFO 5 = 310505-90610 6 = 219895 8 Cos